40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog, Page 36 of 37  Not logged in ELOG logo
ID Date Author Type Category Subject
  154   Sat Jul 17 13:41:45 2010 JanMiscSeismometryRanger

Just wanted to mention that the Ranger is reassembled. It was straight-forward except for the fact that the Ranger did not work when we put the pieces together the first time. The last (important) screws that you turn fasten the copper rings to the mass (at bottom and top). We observed a nice oscillation of the mass around equilibrium, but only before the very last screw was fixed. Since the copper rings are for horizontal alignment of the mass, I guess what happens is that the mass drifts a little towards the walls of the Ranger while turning the screws. Eventually the mass touches the walls. You can fix this problem since the two copper rings are not perfectly identical in shape, and/or they are not perfectly circular. So I just changed the orientation of one copper ring and the mass kept oscillating nicely when all screws were fastened.

  153   Thu Jul 8 16:41:29 2010 JanMiscSeismometryRanger Pics

This Ranger is now lying in pieces in West Bridge:


 DSC02485.JPG
DSC02486.JPG

DSC02487.JPG

First we removed the lid. You can see some isolation cylinder around the central metal 
part. This cylinder can be taken out of the dark metal enclosure together with the interior 
of the Ranger.

DSC02489.JPGDSC02491.JPG

Magnets are fastened all around the isolation cylinder. One of the magnets was missing 
(purposefully?). The magnets are oriented such that a repelling force is formed between 
these magnets and the suspended mass. The purpose of the magnets is to decrease the 
resonance frequency of the suspended mass.

DSC02492.JPGDSC02493.JPG
DSC02496.JPGDSC02494.JPG

Next, we opened the bottom of the cylinder. You can now see the suspended mass. 
On some of the following pictures you can also find a copper ring (flexure) that was 
partially screwed to the mass and partially to the cylinder. Another flexure ring is 
screwed to the top of the test mass. I guess that the rings are used to fix the horizontal 
position of the mass without producing a significant force in vertical direction. The 
bottom part also has the calibration coil.

DSC02502.JPGDSC02505.JPG

Desoldering the wires from the calibration coil, we could completely remove the mass 
from the isolation cylinder. We then found how the mass is suspended, the readout 
coil, etc.:

DSC02509.JPG DSC02513.JPG
DSC02514.JPGDSC02516.JPG

  152   Wed Jul 7 00:14:55 2010 Vladimir DergachevNoise HuntingSUS2e-9 rad/sqrt(Hz)
Here is a fresh plot:

https://ldas-jobs.ligo.caltech.edu/~volodya/tiltmeter/tw_driver3/preamp4/fine_tilt_spectrum_zoomed.png

Changes: reduced noise in preamps, reduced gain in fine channels so that
they just reach 1e-10 rad/sqrt(Hz) - this produces large overlap without
having to trim LVDTs.

We are now under 2e-9 rad/sqrt(Hz), but there is still work to be done as
can be clearly seen from the correllation plot:

https://ldas-jobs.ligo.caltech.edu/~volodya/tiltmeter/tw_driver3/preamp4/LVDT1_LVDT2_fine_corr.png

Or from overlay of coarse and fine channels:

https://ldas-jobs.ligo.caltech.edu/~volodya/tiltmeter/tw_driver3/preamp4/lvdt1_combined_spectrum_zoomed.png
https://ldas-jobs.ligo.caltech.edu/~volodya/tiltmeter/tw_driver3/preamp4/lvdt2_combined_spectrum_zoomed.png

The negative correllation values are clearly electronic (likely excitation
coil driver).

Some of it subtracts:

https://ldas-jobs.ligo.caltech.edu/~volodya/tiltmeter/tw_driver3/preamp4/fine_combined_spectrum.png

but there is a lot of room until we reach the pink line (which is the
limit of readout electronics for fine channels).
  151   Sat Jun 26 13:59:57 2010 Vladimir DergachevMiscSUSovernight tiltmeter plots
And here are the plots from overnight run:

https://ldas-jobs.ligo.caltech.edu/~volodya/tiltmeter/tw_driver3/preamp1b/

Notes:

* I picked a nice drift segment out of the whole run which showed some
junks in the beginning and a few near the end, possibly caused by external
effects.

* The drift with oscillations is still there. It is likely they are
mechanical:

https://ldas-jobs.ligo.caltech.edu/~volodya/tiltmeter/tw_driver3/preamp1b/coarse_tilt_vs_time.png

* The best spectrum comes from lvdt2. It is likely that LVDT1 receives
extra noise from clamping zeners in the fine channel of its preamplifier.

https://ldas-jobs.ligo.caltech.edu/~volodya/tiltmeter/tw_driver3/preamp1b/lvdt2_combined_spectrum_zoomed.png

* The full spectrum reaches the limit of the coarse channel only at
high frequencies:

https://ldas-jobs.ligo.caltech.edu/~volodya/tiltmeter/tw_driver3/preamp1b/lvdt2_combined_spectrum.png

The fine channel of LVDT2 dips a little bit lower, which is easier to
see on linear X scale:

https://ldas-jobs.ligo.caltech.edu/~volodya/tiltmeter/tw_driver3/preamp1b/lvdt2_combined_spectrum_linear.png

The electronics operates around 6.6 kHz and the conversion to DC is
done digitally. Thus we should see flat noise floor from it, except for
the effect of voltage references which are used both in ADCs and in the
triangle driver and any other noise source that affects the amplitude
(such as a current setting resistor in the triangle driver).

Riccardo - I think it would help to isolate the effects of mechanics
noise from driving electronics if we had a test fixture for LVDTs.
Something like a U bracket for the excitation part of the LVDT and a
screw-on cover for the pickup coil. It would be nice to have a choice
between a plain cover, a cover with a slot and a PEEK cover.

This essentially follows the suggestion Eric made at the last meeting,
except I would avoid usage of all-analog readout as I am not confident I
can debug it easily. We can still do it as a confirmation once we know
what the baseline curves are from our current system.
  150   Thu Feb 18 17:43:15 2010 MottMiscSUSPiezos

 

I finished crimping all the connectors we will need for the piezos.  We are now just waiting for the new structures to arrive so we can start gluing the piezos on.

  149   Fri Feb 12 11:35:19 2010 MottMiscSUSPiezos

Engineering was very helpful showing me how to make the leads we need for the piezos; I will go crimp some more at the beginning of next week. 

The new structures should be coming in soon, so we will have a dedicated structure for the piezo damping, at which point we can really get cracking.

  148   Fri Feb 12 11:33:04 2010 MottMiscSUSState of the Shaker

The Shaker project is coming along nicely.  I am currently looking into using the built-in ability to download a waveform to the front end to do the sweeps, but we are running into memory problems, and I get the sense from Tony that it was not really designed to do this.  Currently we are able to download a waveform to the frontend, run the generator according to it, and make a measurement over a full run of the waveform.  If we can crack the limited time constraint and figure out the averaging, this is going to be the most straightforward solution.

I am working, in parallel, with Gert (who is out of the office at the moment) on using pure script to do this, although I am worried about starting and stopping the generator so frequently.  Apart from anything else, there is a slight hang in the frontend when the generator start method is called; it is not noticeable when the button is pushed in the app, but I think it adds quite a bit of latency to the program.  I am still waiting to hear from Gert about how to acquire a time series; hopefully we can figure that out by early next week, since it is critical.  I am also not entirely sure how we force the program to do all the analysis on the time series after it is acquired.  Ideally we would want the analysis to run in parallel and update the frontend continuously, but I am not sure this is possible with VBA (I don't think you can do multithreaded programming) and I am not sure I would know how to do so even if it is!

 

 

  147   Tue Jan 26 22:22:17 2010 ZachMiscSUSAOSEM LED, PD current comparison

Norna and Rich: I am sorry for taking so long to get you this measurement. I plan to do the noise measurements on the standalone LEDs this week.

The following table gives the current through each OSEM's LED (measured using the voltage drop across the 238-ohm resistor in series), as well as the measured photocurrent (the DC output of the amplifier divided by its transimpedance gain of 100,000 V/A), and the ratio of the two. The plot from the previous post is reproduced below for analysis--I realized that I did still have this plot saved in units of V/rHz. In some cases (e.g. #3, #4), the noise level seems to be correlated to the photocurrent, but not all of them follow this pattern. The issue of #1 being significantly quieter than the other set remains, as well.

AOSEM LED and PD Current Comparison
AOSEM # I_LED (mA) I_PD (uA) I_LED/I_PD
1 35.1 31.4 1120
2 35.5 38.5 923
3 35.5 57.6 616
4 35.4 30.5 1160
5 35.4 42.9 826
6 35.5 48.2 737
7 35.5 39.2 906
8 35.4 39.8

889

9 35.2 34.0 1040
10 35.5 43.8 810

 

AOSEM_comparison_1_12_10.png

 

  146   Wed Jan 13 14:13:03 2010 ZachMiscSUSASOSEM comparison

EDIT: I have calibrated the y axis of the plot to meters

Last night, I got around to testing some of the other AOSEM samples, to see how the noise varied between them. What I found was rather strange: the noise in all the new ones (#s 2-6) was about the same, but they were all quite a bit noisier than the previous one I have been testing (#1). The only difference between them, as far as I can tell, is that the first specimen has a coil wound around it already, while the others just have a rubber band. Also, the newer ones all have an impedance of ~ 44-45 ohms, while I measured 47 ohms for the first (though, among the new ones, the slight variation in Z seems to have no correlation with the small differences in noise level). For those wondering, YES, I did remeasure the noise in the 1st one; I am not using old data.

Either my meddling with the old one has somehow made it quieter or something is amiss.

AOSEM_comparison_m_1_12_10.png

  145   Wed Jan 6 14:43:38 2010 ZachElectronicsGeneralWaitDone Error ??

Same thing happened when I went in this morning. I left it for a while (on this time) and it seems to be working. We might want to have this machine looked at.

Quote:

 The SR785 I am using for the AOSEM noise measurements (the one that was in the TCS lab) doesn't seem to want to boot up all the way. After sitting at the "Backup OK" screen for ~30 secs, a dialog box pops up reading "WaitDone Error", after which the machine reboots. This continues forever.

I remember this having happened when I first liberated it from the TCS lab a few weeks ago. I turned the thing off for a while and it eventually worked just fine. I tried the same thing now and it didn't work. I am going to give it another go in the morning.

Has anyone experienced this error before?

 

  144   Tue Jan 5 21:14:46 2010 ZachElectronicsGeneralWaitDone Error ??

 The SR785 I am using for the AOSEM noise measurements (the one that was in the TCS lab) doesn't seem to want to boot up all the way. After sitting at the "Backup OK" screen for ~30 secs, a dialog box pops up reading "WaitDone Error", after which the machine reboots. This continues forever.

I remember this having happened when I first liberated it from the TCS lab a few weeks ago. I turned the thing off for a while and it eventually worked just fine. I tried the same thing now and it didn't work. I am going to give it another go in the morning.

Has anyone experienced this error before?

  143   Fri Dec 18 11:56:12 2009 ZachMiscSUSNew AOSEM LED problem

 I picked up the other AOSEM from the 40m today, so that I could compare it with the one I've measured in an attempt to get to the bottom of the noisy LED problem. It began uneventfully: I measured the impedance of the LED by connecting an ammeter in series and slowly increasing the voltage. I got ILED = 36 mA at about VLED = 1.7, giving ZLED ~ 47 ohms.

Then, I powered up the LED driver, and tested it with a 50-ohm resistor (as usual), measuring V ~ 1.7 across it. Having confirmed that everything was working properly, I hooked the LED up, and measured NO current. I hooked directly back up to the DC supply and found the same result. The thing appears to be blown, but I have no idea how. I went through every precaution I have been taking with the other OSEM, which worked fine when I switched it back in. Crap.

One thing I noticed before I hooked anything up was that the small white pieces attached to the LED and PD on either side of the OSEM opening were very loose when compared to the other OSEM. When I first measured no current, I tried applying some pressure to the LED side, and some current flowed across, but only about 1/10 of what it should have been.

  142   Thu Dec 17 10:50:16 2009 ZachMiscSUSAOSEM noise measurement

 I retook the measurement from the previous post, since the noise in the amp appeared much higher than it should. It looks much better now, but still not great. Above around 3 Hz, the amp noise is at the LISO-predicted level of ~8e-12 m/rHz equivalent displacement noise. Below this, it seems to show some 1/f-ish noise back to DC. LISO predicts some low-frequency noise as well, due to the increase in the AD743's current and voltage noise there, but it doesn't begin until below 1 Hz, and it doesn't seem quite as steep. I'm trying to figure out what is causing this, but the best solution might be to move to a more solid, soldered arrangement.
Details:
LED: Voltage across the LED (DC value 1.61 V), AC coupled
Bright: Noise at output of amp, LED on and PD connected, DC coupled with bandpass
Dark: Noise at output of amp, LED off and PD connected, DC coupled with bandpass
Amp: Noise at output of amp, PD disconnected, AC coupled (I checked this DC coupled and through the bandpass, as well, but there is no significant difference)
Amp (theoretical): LISO-predicted noise at the output of the amp with no input
AOSEM_noise_plot_12_15_09.png

 

  141   Tue Dec 15 10:58:56 2009 ZachMiscSUSPD front end noise

10k was a typo--fixed.

Quote:

This is the LISO estimate of the PD front end noise. It could be improved somewhat by using a higher value resistor, but there's no point.

The shot noise level for the 20-40 uA of current we have is more than 1 pA/rHz so we should be OK above 30 mHz.

So even the level of the dark noise below seems too high and also the 10,000 V/A statement. The feedback resistor we had used to be 100k...

 

  140   Tue Dec 15 09:28:39 2009 ranaMiscSUSPD front end noise

This is the LISO estimate of the PD front end noise. It could be improved somewhat by using a higher value resistor, but there's no point.

The shot noise level for the 20-40 uA of current we have is more than 1 pA/rHz so we should be OK above 30 mHz.

So even the level of the dark noise below seems too high and also the 10,000 V/A statement. The feedback resistor we had used to be 100k...

Attachment 1: dcpd.png
dcpd.png
  139   Tue Dec 15 01:48:22 2009 ZachMiscSUSAOSEM noise measurement

 I tracked down some more AD743s at Wilson House 2.0 today (thanks to Rich). I was then able to simultaneously use 743s to both drive the LED and amplify the readout. Below is an 800-line DC - 25 Hz noise spectrum of

- The voltage across the LED (DC level: VLED = 1.59 V) -- AC coupled

- The output of the amp with a fully lit PD (DC level: Vout,full = -2.95 V) -- DC coupled, through bandpass filter

- The output of the amp with the LED out -- DC coupled, through bandpass filter

- The output of the amp with an open input -- DC coupled, through bandpass filter

all calibrated to equivalent displacement noise. For the LED plot, this was done by using the measured current ratio between the LED and the PD when fully lit along with the measured OSEM response of 0.05 A/m. The other three were converted using this response along with the transimpedance gain of the amp (100,000 V/A). For all measurements, the OSEM was covered by a box, and the circuit was draped with a cloth and put under a box within another box (to reduce air currents).

AOSEM_noise_12_14_09.png

The funny low-frequency junk from the previous driver spectrum is gone--thanks to the isolation from air currents--but the line seems a bit higher overall (trying to figure out why). There also still seems to be the funny effect of noise added by the LED above the level of the voltage noise across it, and I think it's somewhat strange that the "Dark" noise is lower than the amp noise with no input. We can probably still do better..

  138   Tue Dec 8 10:07:39 2009 ZachElectronicsSUSSallen-Key filter attenuation limit

 Last night, I was going to retake the noise measurement with the added elements that Rana suggested, but instead I spent a ridiculous amount of time trying to figure out what is going on with my Sallen-Key filter. I now know a lot more about their limitations, but am still at a loss as to what is happening in this case. The problem is that no one seems to be using a corner frequency anywhere near this low (at least not anyone trying to explain these filters). The following is a comparison of an ideal 2nd-order Butterworth filter and a real one using an LT1464.

butterworth_limitation.png

The filter behaves as expected a ways past the corner frequency, rolling off at 40 dB/decade. As the signal increases in frequency, the capacitors' impedances decrease, until (at point 'b'), they fall below the output impedance of the amp, causing the response to climb at 40 dB/decade. This happens for a short while, until the unity-gain bandwidth frequency (~ 1-10 MHz, 4 MHz for the AD743) of the op amp is reached, and the filter can attenuate no more ('c'), so the response flattens out to 0 dB/decade.

Different component values affect the high-frequency behavior of the filter, as shown below.

S-K_different_components.png

 

This makes sense, since with smaller capacitors it takes a higher frequency to fall below the output impedance of the amp. In any case, though, the final flatline always happens near the UGB frequency. The following is a plot of the transfer functions I measured (also in last post). I did not change the C values--only the R's--so the corner frequency is different in each case. What I observe is some R-dependent maximum attenuation, which sets on well before the 4-MHz UGB frequency of the AD743. The strangest part is that for small enough R this maximum "attenuation" is actually a positive gain.

S-K_respons_vs_R.png

I suppose it is not a huge deal, as I can increase the stopband attenuation as R while only increasing the Johnson noise as 2 * sqrt(R), but it would be nice to get some insight into what 's going on.

NOTE: While it appears that the high-frequency flatline in the other plots occurs for a different reason, it still seems to be R-dependent. I could not find any explanation for what determines this asymptotic behavior.

  137   Mon Dec 7 12:01:14 2009 ranaMiscSUSAOSEM LED Driver noise (with AD743)

The low frequency noise looks pretty good now. The funny shape is most likely a thermal transient due to having not enough insulation. You need to droop some Kleenex over the circuit to stop the thermal air currents and then put a second box over the first box. Then its probably best to sit outside of the room when taking the measurement to reduce the human noise.

  136   Mon Dec 7 02:59:38 2009 ZachMiscSUSLED Driver noise (with AD743)

 I retook the measurement from my last post, this time using an AD743 in place of the OP27 (per Rana's comment). The results are below.

Low-frequency:

noise_comp_low.png

RA: Although it says m/rHz, this is not measured displacement noise, but rather estimated displacement noise due to the LED noise. The previously measured conversion from the LED RIN to apparent displacement is used to convert from the voltage noise of the LED driver to the contribution to the OSEM's displacement readout.

High-frequency:

noise_comp_high.png

Seems better than before, but not quite what expected. I observed that the transfer function of the S-K filter was what it should be up until a decade or two above the corner frequency, after which it appeared to spring zeroes out of nowhere and level off at high frequency. I tried to see what would happen if I changed the resistor values, and the following plot is what I got.

S-K_respons_vs_R.png

This plot seems familiar from my electronics courses, but I haven't put a finger on what is causing this behavior yet. I'm sure that the answer is somewhere in H&H (or in the brain of a kind soul who happens to be reading this--wink wink).

  135   Thu Dec 3 02:35:39 2009 ranaMiscSUSLow-noise LED driver

 

 OP27 current noise is too high - use AD743.

  134   Wed Dec 2 18:29:07 2009 ZachMiscSUSLow-noise LED driver

 Yesterday, I rebuilt Rana's low-noise LED driver in the Bridge elab. It is based on a 2nd-order active lowpass filter (using the Sallen-Key topology). The schematic is shown below. The circuit is essentially the same as the one Rana posted a few days ago, only the R and C values are all around twice what they are in his schematic. This results in the same corner frequency of fc = 0.03 Hz.

 

schematic.png

 

I hooked it up and measured Vsk,out = 9.76 V. I then used it to drive a 50-ohm resistor, and measured VLED = 1.71 V, then measured the current to be ILED = 33.5 mA.

After ensuring that it was supplying the correct voltage, I hooked it up to the LED and took a spectrum of the voltage noise across it over the two frequency bands I have been using in previous posts. The following are comparison plots of the noise here and the noise with the simple RC filter used before, calibrated to displacement noise.

Low-freq:

noise_comp_low.png

RA: although these plots have Displacement in the y-axis, they are NOT measurements of actual displacement noise. They are estimates for the contribution to the displacement noise made by the LED RIN based on measurements of the voltage noise across the LED.

High-freq:

noise_comp_high.png

Something is clearly wrong: not only is the new configuration worse at lower frequencies, but the rolloff seems to go as 1/f and not 1/f2. Investigating after dinner...

 

  133   Tue Dec 1 00:49:19 2009 KojiElectronicsSUSLED Driver circuit

The situation is well illustrated in the following application note of Analog Devices:

Low Noise Amplifier Selection Guide for Optimal Noise Performance

Even though the graph is created for 1kHz, it is very clear that AD743
has superb performance for high source impedance purposes:
combination of low current noise and low voltage noise.

If the source impedance of Rana's circuit (400kOhm@DC) are reduced to 20K or so,
OP-27 type OPamp can come into the scope. However this means we need to use 10 times
larger capacitors. This is almost impossible for now, though innovation on the caps
can change the situation.

LT1028 is an AD797 type opamp. This works greatly with the smaller source impedance.

Attachment 1: AN_940_Page_08.png
AN_940_Page_08.png
  132   Mon Nov 30 18:53:28 2009 ZachElectronicsSUSLED Driver circuit

Not having much luck. I found the LT1028, which has 10x better low-frequency voltage noise (35 nVpp, 0.1 Hz to 10 Hz), but its current noise is worse by a ridiculous factor of 1000:

Screen_shot_2009-11-30_at_6.51.02_PM.png           Screen_shot_2009-11-30_at_6.51.16_PM.png

 

 

Quote:

I found a Quad Opamp OP497 (neither dual nor single!), but this is not enough to expel AD743.
Dis-continuation of OP497 is also close except for one SMD package.

AD743 (reference):
LF Voltage Noise: 0.38 V pp, 0.1 Hz to 10 Hz
Voltage Noise: 2.9 nV/√Hz @ 10 kHz
Current Noise: 6.9 fA/√Hz @ 1 kHz

OP497:
LF Voltage Noise: 0.3 V pp, 0.1 Hz to 10 Hz
Voltage Noise: 15 nV/√Hz @ 1kHz
Current Noise: 5 fA/√Hz @ 1kHz

Quote:

What we really need to make this good is a part with just as good of an input current noise spec as the AD743 but 3x less voltage noise at 0.1 Hz. I offer one cookie to whomever can find an opamp that fits those parameters.

 

Attachment 1: Screen_shot_2009-11-30_at_6.51.02_PM.png
Screen_shot_2009-11-30_at_6.51.02_PM.png
  131   Thu Nov 26 15:17:01 2009 KojiElectronicsSUSLED Driver circuit

I found a Quad Opamp OP497 (neither dual nor single!), but this is not enough to expel AD743.
Dis-continuation of OP497 is also close except for one SMD package.

AD743 (reference):
LF Voltage Noise: 0.38 V pp, 0.1 Hz to 10 Hz
Voltage Noise: 2.9 nV/√Hz @ 10 kHz
Current Noise: 6.9 fA/√Hz @ 1 kHz

OP497:
LF Voltage Noise: 0.3 V pp, 0.1 Hz to 10 Hz
Voltage Noise: 15 nV/√Hz @ 1kHz
Current Noise: 5 fA/√Hz @ 1kHz

Quote:

What we really need to make this good is a part with just as good of an input current noise spec as the AD743 but 3x less voltage noise at 0.1 Hz. I offer one cookie to whomever can find an opamp that fits those parameters.

  130   Thu Nov 26 02:59:35 2009 ranaElectronicsSUSLED Driver circuit

We want to have a simple low noise circuit to drive the LED. Our plan is to use the AD587 followed by a filter/buffer.

Requirements:

from 0.1-10 Hz, produce less RIN in the LED light than shot noise by a factor of 3.

With 35 mA of LED drive, we get ~35 uA of photocurrent (no magnet/flag). The shot noise of 35 uA is ~3.5 pA/rHz.

So the RIN from shot noise is 1e-7. So we shoot for a RIN of 3e-8 from the LED.

 

The AD587 voltage reference has a relative noise of 1e-7 at 0.1 under very good conditions (perhaps our vacuum system will be so kind). So we have to get a factor of 3 filtration at 0.1 Hz.

The following circuit should it for us: its a 2nd order Butterworth implemented in a Sallen-Key configuration. The noise is reasonable and the cutoff frequency is so low (0.03 Hz) because of the latest in capacitor technology.

We can buy metal poly caps which are as large as 47uF and have a reasonable physical size and tolerance and noise.

On page 2 of the plot you can see that the noise performance of this filter is limited by the input voltage noise of the FET opamp (op1) (AD743 - soon to be obsolete). The noise of the BUF634 (op2) is insignificant in this configuration. What we really need to make this good is a part with just as good of an input current noise spec as the AD743 but 3x less voltage noise at 0.1 Hz. I offer one cookie to whomever can find an opamp that fits those parameters.

These images show the circuit diagram (left) and the proto setup (right):

IMG_0232.JPGIMG_0226.JPG

update: added a 230 Ohm series resistor between the BUF634 output and the LED to step the voltage down to the 1.7V that the LED wants.

Attachment 3: sallenkey2.pdf
sallenkey2.pdf sallenkey2.pdf
  129   Tue Nov 24 23:17:02 2009 ZachMiscSUSAD587KN voltage noise

I got a few AD587KN (high-precision 10V reference) samples today from AD. I hooked them up to see how much quieter my DC supply would be. The results are pretty good, with the voltage noise reduced by a factor of 5-10 throughout. The first two attachments below are comparisons of the noise in

1. The +12V regulator (MC78M12) alone

2. The AD785KN reference with V_in = +12 V provided by the regulator

3. The same as in 2, only now with an additional "noise reduction" capacitor (a 1-uF capacitor from pin 8 to ground forms a LPF with an internal 4-k resistor, giving a corner frequency of 40 Hz to reduce high-frequency noise),

plotted with the same frequency ranges and settings as those in the previous post.

The reference comes very close to its noise spec of 100 nV/rt(Hz) @ 100 Hz. The only issue is that it seems to have much more line pickup than the regulator (which seems almost completely insensitive to line noise), and this is worsened by the extra capacitor. Attachment 3 is a close-up of the low-frequency spectrum around 60 Hz. I suspect that this will be alleviated somewhat when I move away from the breadboard phase.

I want to rig this up so that I can stabilize the supply voltage to the transimpedance amp and LED, but in order to do so I will need to build a higher-current source using a power transistor, like either of those shown in attachment 4 (the AD587LN is only able to provide <10mA). 

Attachment 1: ref_noise_low.png
ref_noise_low.png
Attachment 2: ref_noise_high.png
ref_noise_high.png
Attachment 3: line_noise.png
line_noise.png
Attachment 4: high_current.png
high_current.png
  128   Tue Nov 24 12:28:52 2009 ZachMiscSUSAOSEM measurement update

  I was able to take some better measurements last night. I took data in two bands: 0-100 Hz, 0-1.6 kHz, each with 800 lines. This gives us a decent idea of what's going on at low and high frequency. Attached are four plots, two from each band. All measurements were taken with a box over both the OSEM and the readout circuit and the lights out.

The first two are low- and high- frequency comparisons of the noise in the full (bright) configuration as measured with no BPF and AC coupling vs with the BPF and DC coupling. There appears to be no difference apart from the expected effect above the pole at 1 kHz.

The next two are plots of the noise in various components and the full scheme calibrated into equivalent displacement noise. Everything is below ~10e-10 m/rt(Hz) with the exception of line peaks, and again it would appear that we are limited by our measurement equipment.

Some notes:

- The "dark" noise seems to be coincident with the "amp" noise with the exception of some extra pickup that increases at high frequency (seems to be line-related).

- The "LED" noise is coincident with the "supply" noise up until its 8-Hz corner frequency, after which it falls off as expected until it hits an apparent floor around 100 Hz.

- The "bright" noise seems to be coincident with the "supply" noise, while the "dark" and "amp" are much lower. This could be because the supply noise only shows up when there is an appreciable voltage at the output of the amp.

 

Have to think about this for a bit, but the next logical step is to turn the measurement setup into something solid (i.e. soldering, enclosure, etc.).

Attachment 1: BPF_low.png
BPF_low.png
Attachment 2: BPF_high.png
BPF_high.png
Attachment 3: OSEM_low.png
OSEM_low.png
Attachment 4: OSEM_high.png
OSEM_high.png
  127   Sun Nov 22 16:02:09 2009 ZachMiscSUSAOSEM noise measurement

On Friday, Rana and I discovered that my transimpedance amp was oscillating like whoa at about 100 kHz. A little research showed this to be due to the input capacitance of the AD743 (~20 pF). To fix this, I put a 20-pF cap in parallel with the 100k feedback resistor, and that seemed to do the trick.

 

The relevant circuitry is shown in attachment 1. +/- 12 V DC was provided by voltage regulators (7912, 78M12). The voltage across the LED was measured to be V_LED = 1.61 V, and the current through it was I_LED = 31.6 mA, giving Z_LED = 50.9 ohm. The voltage out of the amp with a fully lit PD was V_out = -2.83 V, giving a photocurrent of I_ph = 28.3 uA. 

 

I was concerned about noise that might be imposed by the bandpass filter, so I compared spectra I took with and without it (that is, AC coupled, no BPF, and DC coupled, with BPF). This comparison is shown in attachment 2. There appears to be no difference apart from the aliasing effects at low frequency.

 

After this, I took the real measurement, extending the range to 800 Hz, averaging 100x and with a linewidth of 1 Hz (I realize now that I should probably have done this with a smaller linewidth, so that I could see below ~1 Hz. I will repeat the measurement this week with better low-frequency resolution). The result can be seen in attachment 3, calibrated to displacement noise in m/rt(Hz) using the measured 0.05-A/m response of the OSEM in the linear region. The four lines are:

 

- Bright: noise in the OSEM with a fully lit PD

- Dark: noise in the OSEM with the LED off

- Amp: noise in the transimpedance amp with the input terminated

- V_LED: noise in the LED voltage

 

The first three spectra were taken at the output of the amplifier and calibrated back to meters using the transimpedance gain and OSEM response in A/m. The last was taken across the LED, and calibrated into meters using the values given in paragraph 2. All measurements were taken with the OSEM under a box and with the lights out.

 

It appears that we are still limited by our setup. The "Dark" noise is coincident with the amplifier noise, while the "Bright" noise is coincident with the LED noise. That said, it is fairly comforting that all this noise is at the level of around 10^-10 m or less, as we can probably expect the true noise of the OSEM to be lower than this. We will know this for sure once we have a truly quiet setup (starting with ultra-low-noise voltage references).

Attachment 1: schematic.png
schematic.png
Attachment 2: BPF_noise_comparison.png
BPF_noise_comparison.png
Attachment 3: 0-800.png
0-800.png
  126   Fri Nov 20 06:18:51 2009 ranaMiscSUSbandpass filter

seems OK, as long as the current noise doesn't get you. To make sure you have to terminate the input to this filter and then look at the resulting noise in the SR785.

  125   Thu Nov 19 12:00:04 2009 ZachMiscSUSbandpass filter

 Attached is the transfer function for the bandpass filter I built for the AOSEM readout. The schematic is primitively outlined below. The corner frequencies are what they should be (HP: 100 mHz, LP: 1 kHz)

The Johnson noise for the 1 k resistor is roughly 4 nV/rt(Hz). Frank says that it doesn't make sense to use much lower than 1 k if I'm putting it into an SR785.

 

 

           10 uF                  1 k

      -------| |-------------/\/\/\/\/\/------------------

                        | |

        < |

         160 k  >        _|_  160 nF

        <        ---

        > |

        | |

                V                 V

 

Attachment 1: BPF_bode_11_18_09.png
BPF_bode_11_18_09.png
  124   Thu Nov 19 03:41:12 2009 ZachMiscSUSAOSEM calibration

 Tonight, I calibrated the AOSEM's response in [A/m]. I used a sophisticated rig consisting of:

 

1. One of those anodized Faraday isolator mounts to hold the OSEM

 

2. A translation stage with a screw gauge to jam something into it (gracefully)

 

3. Some DC power supplies and multimeters

 

4. My simple transimpedance amplifier sketched in the previous post (I did not bother upgrading the readout circuit for this measurement since I was just taking a relatively rough DC measurement)

 

---------------

 

I used a 9/64 hex key to simulate the shadowmaking magnet. A picture of the setup is attached below.

 

Some pertinent info:

 

- The current through the LED was maintained at 35 mA throughout the measurement. The measured voltage across it was 1.62 V, giving Z_LED = 46.3 ohm.

 

- The op amp supply voltage was +/- 10 V, and the PD bias was +10 V.

 

- The output voltage of the amplifier with the PD fully lit was 3.04 V (measured before and after the test). Note that this voltage increases slightly as the key is inserted due to reflections. 

 

 

The second attachment is a plot of the photocurrent versus the position of the key (the x axis is shifted such that the key is roughly centered at x = 0, and x < 0 corresponds to the key being further inside). The response of the OSEM in the linear region is roughly 0.05 A/m.

Attachment 1: 100_0362.JPG
100_0362.JPG
Attachment 2: calibration_plot_11_18_09.png
calibration_plot_11_18_09.png
  123   Tue Nov 17 21:23:08 2009 KojiMiscSUSAOSEM test progress

We have LT1021-7 at the 40m, next to the Alberto's desk. This is the VREF for 7V.

Quote:

1. Voltage regulator/reference

Rana thinks that the AD587LN is a good choice of reference given its performance on some LISA tests. I am in contact with AD, and there is no longer a 'LN' package, but I am trying to get samples of the currently manufactured one that is most similar (AD587KNZ).

In the meantime, I am going to find some simple regulators downstairs or at the 40m. 

 

  122   Tue Nov 17 10:33:37 2009 ZachMiscSUSAOSEM test progress

 It's dusty in here...

--------------------------------

 

I was recently commissioned to do some noise measurements on the new  AOSEMS. I set up a humble experiment in the LIGO e-lab to do some preliminary measurements:

 

I made a simple current-to-voltage converter out of an OP27E (using a 100-kohm feedback resistor) to use as the transimpedance amplifier for the readout. This results in a transimpedance of 0.1 V / uA. A simple schematic of the important elements is attached below.

 

DC power was provided without regulators directly from the laboratory DC supply in the lab. The value of 1.7 V across the LED was set such that the current through it was ~35 mA.

 

Rana and I took a few important PSDs (one of the DC supply, one of the OP27E with no supplied current, and two with the setup fully connected--one each with and without the PD covered), all from 250 mHz - 200 Hz, AC coupled. Using a sophisticated estimation method (called, by some, the "pick two points and approximate with a power law" method for lack of something fittingly elegant), we obtained a rough estimate of these spectral densities in order to compare them.

 

These were all converted into equivalent PD current noise. For all but the "supply" noise, this was done trivially by dividing by the transimpedance of the OP27E. For "supply", LED voltage noise had to be converted to PD current noise in the following way:

 

Z_LED = 1.7 V / 35 mA ~ 50 ohm

 

equivalent PD current noise = (I_PD / I_LED) * (measured supply voltage noise / 50 ohm)

 

where the PD-LED current ratio was found empirically to be (I_PD / I_LED) ~ 1 / 1000 by measuring the voltage out of the amp with full brightness (i.e. I_LED = 35 mA, no obstruction) and dividing by the transimpedance (see 2nd figure).

 

The third figure below is a plot of these spectral densities in common units. Somewhat expectedly, the noise of the "dark" configuration seems limited by the supply noise. However, the "bright" line seems to be dominated by something else. I'm not sure I see how it could be anything but the LED itself, but it is worthwhile to repeat this "test" with a better setup.

 

On the to-do list:

 

1. Voltage regulator/reference

Rana thinks that the AD587LN is a good choice of reference given its performance on some LISA tests. I am in contact with AD, and there is no longer a 'LN' package, but I am trying to get samples of the currently manufactured one that is most similar (AD587KNZ).

In the meantime, I am going to find some simple regulators downstairs or at the 40m.

 

2. Bandpass filter

I was advised that it is a good idea to build your own high/bandpass filter instead of relying on the spectrum analyzer's AC coupling function. I will be doing just this.

 

3. Switch to a better op amp

        Like the AD743

 

4. Calibration

I need to find a good way to hold the OSEM in place while I stick something in there with a micron drive without it being unreliably shaky.

 

 

 

 

 

 

Attachment 1: schematic.jpg
schematic.jpg
Attachment 2: 2009-11-12_17.34.58.jpg
2009-11-12_17.34.58.jpg
Attachment 3: noise_comparison.png
noise_comparison.png
  121   Thu Feb 28 13:46:16 2008 David YMMiscOMCTip Tilts - Susp. Wire shipped
As per Adam Mullavey's request, I have shipped the remainder of the (cleaned) suspension wire for the tip tilt assemblies to LLO for the out of vacuum tip tilts. Unclean the wire as necessary, as Bob says the bake job is trivial for it.
-Dmass
  120   Fri Jan 25 22:08:45 2008 waldmanLaserOMCOMC ready for shipment
The OMC has been boxed in its spiffy black box and is ready for shipment.

We mounted the OMC in its frame, removing any loose mechanical parts and tightening up the nylon tipped set screws. Then we swaddled each optic in tin foil and lint free cloth before wrapping the whole thing in a double layer of the shiny stuff. Finally we wrapped the whole kaboodle in a double layer of ameristat before shoe-horning (literally) it into the box. The package is about 1/4" too long, so it required some effort to get it into the box, and will doubtless require similar effort to remove it.

The OMC in its box was left on top of its shipping container in 056.
  119   Wed Jan 23 22:52:21 2008 Norna RobertsonMiscSUSOMC SUS assembly adjustments made and lessons learnt
Calum and I have been working in the 40 m annex re-assembling and testing the OMC SUS after the parts have been cleaned and baked and various parts modified. We have had several problems which needed to be addressed and made several observations of items which need attention for the next OMC. Also we note several points which need to be followed when the OMC is being prepared at LLO for input into the vacuum chamber. The following are in no particular order.

1) Problem with one of the blades
When we got the double pendulum hanging today we noticed that one of the lower blades was interfering with part of the top mass - hitting a block which holds one of the screws for adjusting the position of the clamp for the upper wire ( for adjusting pitch). The blade was 13S - one of those which had an angled clamp of 3.5 degrees (the larger of the angles - the blades at the other end of the bench have 3 degree angles). These large angles are being used to counteract the fact that the bench weighs more than these blades were designed for. The blades thus have a convex curvature (looking from above) and the crown of the bend was where the interference was occuring. The puzzling bit is why only one blade showed this, and not its partner. We tried another 3.5 degree clamp ( in case it was a fault with the clamp) but saw the same effect. We then tried reducing to 3 degrees for that blade - still almost touching. We then reduced the angle for that blade to 2.5 degrees. This worked Ok - and when the masses were hanging again all the blade tips appeared to be at approximately the same height (measured using a steel rule). So the dynamics should be OK. Puzzle remains why only one blade showed this behaviour. Note that for the LHO OMC we will be getting new blades designed for the mass they will be taking, and so we should not need to use angled clamps as large as for the LLO suspension Hence this interference problem should not arise again.
Lesson for the future - need to characterise all blades with the full range of angled clamps to be used. (We did not have time to do this for this suspension).

2) Test of new method of hanging silica bench (using metal bench)
Yesterday with Ken Mailand's help we tried out the new way of suspending the bench using Ken's bench holder with one of the metal benches and the two lab jacks. We learnt that to get the right range of height adjustment to lift the bench to put the discs on the lower wire clamps and then lower the bench again the mating pieces on the lab jacks had to mate with the central portion of the handles at each end rather than the top part of the handles. Also it looks like the lower EQ stop holder (which sits under the bench) should be put in place before the bench is brought in - otherwise difficult to get into place. Finally we note that this job is a three person job - need one on each end to raise and lower the lab jacks and one to fit the discs and then guide the discs and clamps into the holes and check they are seated correctly as the bench is lowered.

3) Height of bench.
We measured the height of the lower surface of the metal bench above the optics table to be (160.5 - 38.75) = 121.75 mm. Requirement is 101.6 + 20 = 121.6 (+/- 2mm) so we are OK. This was without any extra mass added to the top mass. We note that the tabelcloth is close to lower end of its range, but should be OK. However the final alignment needs to be done when the OSEMs are in place.

3) EQ stops and parts which should be removed before putting into vacuum chamber.
For sending the OMC to LLO we put on double nuts on all the EQ stops. This limits the range of the stops and in particular the ones under the bench were not long enough to reach the bench when double nuts are used. So we had to improvise the position of the EQ holder by putting ~ 1/2 inch "shims" to raise the holder ( the shims were three of the masses made for attaching to the metal bench to increase its mass). We also removed all "loose" parts - the adjustment mechanisms for the top blade yaw positions, the magnets and flags, the set screws for locking the larger vertical stops used to hold the top mass, and the screws used to alter the pitch adjustment clamp positions.
When the OMC is reassembled at LLO and made ready for installation, all the second nuts on the EQ stops should be removed. Also once the pitch adjustment using the moveable clamps is set correctly, the screws for doing those adjustments should be removed again. Basically all loose screws should be removed.

4) Parts needing modifications
i) EQ cross pieces holding the plate which goes over the bench need to be reduced in length to fit between the structure legs.
ii) Blind holes in EQ corner brackets.

5) Point to note for next OMC.
i) The slots in the structure which take the dowel pins for alignment need to be lengthened to allow the tableloth the full range of movement which the slots for the attachment screws would allow.
ii) The targets for aligning OSEMS need a hole in the centre for the flag.

More to follow tomorrow
Also we took lots of pictures today. will put relevant ones into installation document.

OMC is due to be crated tomorrow for pick-up on Friday am.
  118   Wed Dec 26 22:39:52 2007 waldmanLaserOMCDCPD beam dumps
I manufactured 2x 1" x 1" black glass beam dumps by vac-sealing the appropriate sized squares to DCPD tombstones. I then locked the cavity and looked at where the DCPD reflections went. The REFL DCPD is no problem, but the choices for the TRANS DCPD aren't so hot. As shown in the enclosed photos, the three options are ugly. I lean towards option 1, but the other two are possibilities as well.

To be clear,

Option 1: beam dump close to the PD at an angle to the beam splitter. Spot is centered on the dump and a little of the tombstone hangs off the edge

Option 2: beam dump close to the PD, parallel to the beam splitter. In order to get the spot on the black glass, the tombstone is very close to the beamsplitter. (This option might be more attractive if the tombstone is rotated 180 degrees.) The tombstone doesn't hang off the edge.

Option 3: We can catch the beam after the beam splitter if we hang the tombstone significantly off the edge, like by half. You'll see in the photos

www.ligo.caltech.edu/~swaldman/BeamdumpPics
  117   Wed Dec 26 15:46:22 2007 waldmanLaserOMCDCPD centering
I used the drilled out radiator with both DCPDs and confirmed that I could center the beams. I also found the reflected beams and confirmed that a) there is clearance for the beamdumps and b) the location of the beams. Note that by some freak of fate, the transmitted DCPD ghost beam is only at about +15mm above the breadboard. I should mount the beamdump a bit low to accomodate it.


Quote:
I can use a CCD camera and a macro lens to see the transmitted cavity beam on the DC photodiodes. I don't know which camera b/c there is no name on it, but it is a square edged camera with a 1/4-20 thread on the mount and has a many pinned cable running to a PS-12SU power supply. I also don't know which macro lens, but tis C-mount, about 4 inches long by 1.5 inches diameter with zoom, focus and f-stop adjustments (all unmarked).

When I look at the face of the transmitted PD, I see the three pins behind the diode chip on the left side. I can move the PD such that the beam is centered on the diode, but this is all the way to one side of the radiator clearance holes. When I maximally misalign the diode, the beam overlaps the diode edge on the side of the feed thru pins.

For the reflected DCPD, I cannot center the beam. Viewing the face of the PD with the two visible feedthru pin son the left side of the diode chip, the beam is low to the left and I cannot center it. I will try again tomorrow with the radiator with the oversize thru holes and see if I can center it. I think it may be a good idea to use an oversized radiator on the transmitted DCPD as well.
  116   Sun Dec 23 18:40:12 2007 waldmanLaserOMCDCPD centering
I can use a CCD camera and a macro lens to see the transmitted cavity beam on the DC photodiodes. I don't know which camera b/c there is no name on it, but it is a square edged camera with a 1/4-20 thread on the mount and has a many pinned cable running to a PS-12SU power supply. I also don't know which macro lens, but tis C-mount, about 4 inches long by 1.5 inches diameter with zoom, focus and f-stop adjustments (all unmarked).

When I look at the face of the transmitted PD, I see the three pins behind the diode chip on the left side. I can move the PD such that the beam is centered on the diode, but this is all the way to one side of the radiator clearance holes. When I maximally misalign the diode, the beam overlaps the diode edge on the side of the feed thru pins.

For the reflected DCPD, I cannot center the beam. Viewing the face of the PD with the two visible feedthru pin son the left side of the diode chip, the beam is low to the left and I cannot center it. I will try again tomorrow with the radiator with the oversize thru holes and see if I can center it. I think it may be a good idea to use an oversized radiator on the transmitted DCPD as well.
  115   Sun Dec 23 17:21:57 2007 waldmanLaserOMCQPD3 centering
In the current cabling scheme, QPD3 is the Far field QPD mounted to the breadboard. With the cavity on a "pretty good alignment" I can physically move the QPD and get the signals to go +/- 1. In other words, I can physically move the QPD to zero its signal, which is the whole purpose of this test. The QPD matrix looks like [1 1 1 1; 1 1 -1 -1; 1 -1 -1 1]; Pitch looks right, yaw might be backwards. Note that the QPD is in a "plus" orientation.

I can center QPD4, the "near field" QPD just fine as well. In this configuration.

One caveat going forward - This alignment is going to be a real pain. In order to be in the linear range of the QPD, we need to be +/- omega/10 or something. In other words, with our 500 micron waist size, we need to get within 50 microns. We should get a translation stage made for this.
  114   Sat Dec 22 18:05:07 2007 waldmanLaserOMCHeater bonding (almost) successful
The heater is now (re)bonded to the OTAS using VacSeal instead of the thermal epoxy. Unlike the expensive high-quality epoxy, VacSeal has supported 4 cycles of connector on/off.

Unfortunately, I didn't bond the heater as straight onto the aluminum as I had hoped. The connector is still about 5 degrees from the vertical so that the connector body sticks up close to the beam (but not as close as it might be. The mode matching is sooooo terrible (because I used an unlabeled lens) that it is impossible to identify the little bit of scattered light coming off the connector. Some of it could be (and probably is), junk light. I believe the closest approach of the connector is ~3 mm, but in the event we measure the finesse to be low, we might want to consider reworking the connector to get it down out of the beam.
  113   Thu Dec 20 19:12:11 2007 waldmanLaserOMCStressful reattachment of heater
Photos may follow eventually, but for now here's the rundown. I scraped the heater clean of the thermal epoxy using a clean razor blade. Then I stuffed a small piece of lint free cloth in the OTAS bore and wrapped the OMC in tin foil. With a vacuum sucking directly from the face of the OTAS, I gently scraped the glue off the OTAS aluminum. I wiped both the OTAS and the heater down with an isoproponal soaked lint-free cloth. I put a thin sheen of VacSeal on the face of the heater, wiping off the excess from the edges with a cloth. Then I clamped the heater to the OTAS using 2" c-clamps from the tombstone back to the heater front, making sure the alignment of the OTAS was correct (connector on the absolute bottom, concentric with the OTAS outer diameter). I added a second clamp, then beaded the outside of the joint with a little bit extra VacSeal, just for kicks. I'll leave it covered at least overnight, and maybe for a day or two.

sam
  112   Thu Dec 20 19:11:40 2007 waldmanComputingOMCHoWTO build front ends
For instance, to build the TPT front end code.

  • Save your file /cvs/cds/advLigo/src/epics/simLink/tpt.mdl
  • go to /cvs/cds/advLIGO on the TPT machine
  • do make clean-tpt tpt install-tpt
  • do rm /cvs/cds/caltech/chans/daq/C2TPT.ini (this step is needed because the DAQ install code isn't quite right at the time of this writing.
  • do make install-daq-tpt
  • run starttpt to restart the tpt computer.

Enjoy.
  111   Thu Dec 20 11:38:46 2007 aivanovComputingOMCframe builder fixed
I put the things back the way they were before we borrowed the ADC board. I can see the TPT data now. Alex
  110   Wed Dec 19 23:10:27 2007 waldmanComputingOMCFramebuilder broken
The framebuilder on OMS isn't doing anything. I can't get data using dataviewer or DTT. Thankfully, I can still use an analog scope to get data.

sam
  109   Wed Dec 19 23:09:43 2007 waldmanLaserOMCOMC relocked
The OMC has been relocked in preparation for final diode alignment, final QPD aligment, and adding the beam blocks. The mode matching is really terrible, so it makes alignment a little difficult because there is a high, high order mode close to the 00 that is making problems for the vertical alignment.

sam
  108   Wed Dec 19 15:29:14 2007 DmassLaserGeneralTip Tilts Being Baked
The Tip Tilt mirrors are now being baked with Bob. They should be ready in time.

The traveler has DCC #E070354-00-X.

As per Bob's know-how, they are cooking at:
80 C for magnets
200 C for steel;
120 C for the Al, Cu, Bronze, and teflon

davidym@caltech.edu
  107   Mon Dec 17 19:39:52 2007 waldmanLaserOMCFiber seems to be broken
The 50 m fiber running from Rana's lab to 056 seems to be broken. I can't get any light through it to save my life. A 5 meter fiber couples like child's play. I think we should acquire a fiber coupler - then I will couple light into the 5 m fiber that works fine and couple it to the 50 m fiber and prove that its broken. Only then will I go pull the installed fiber from the 40m clean room.

sam
  106   Mon Nov 26 16:05:43 2007 Norna RobertsonMiscSUSMass of OMC silca bench (measured 16 Nov 2007)
For the record here is a note sent round OMC-SUS colleagues on 16th November.

--------------------------------

A team of Chris, Chub, Helena and me ably led by Sam (who did the scary lifting) got the OMC bench dismounted from its suspension and weighed today. It is now sitting on its custom plate with dark side down (minus the preamp boxes) in the centre of the optics table in room 56. It will remain there for the forseeable future. The metal parts of the suspension will be disassembled by Chris with Ken Mailand's help starting Monday.

Results of weighing

1) Optical bench including preamps and the wire bundle to the heater, pzt etc = 6.514 kg
2)Preamp cable (mock-up) = 24.4 g (without the connectors to the preamp)
3) The two leaning towers of counterweights (which were situated close to the far corners of the mass - NOT above the holes where they need to be for fixing in place) = 200.2 g and 187.7 g

I remind you all that there are 2 more tombstones ( fused silica 8 mm thick x 30 mm tall x 32? mm wide with 18 mm diam hole- size to be checked from CAD drawing) and two pieces of black glass 1 inch by 1 inch by 3 mm ( I noted at our telecon that these weigh 8.5 gm total) to be added to the bench - situated close to the DC photodiodes which sit on bright side under the preamps.
  105   Thu Nov 15 17:09:37 2007 pkpUpdateOMCVertical Transfer functions with no cables attached.
[Norna Pinkesh]

The cables connecting all the electronics ( DCPDs, QPDs etc) have been removed to test for the vertical transfer function. Now the cables are sitting on the OMC bench and it was realigned.
Attachment 1: VerticaltransferfuncnocablesattachedNov152007.pdf
VerticaltransferfuncnocablesattachedNov152007.pdf VerticaltransferfuncnocablesattachedNov152007.pdf VerticaltransferfuncnocablesattachedNov152007.pdf VerticaltransferfuncnocablesattachedNov152007.pdf
ELOG V3.1.3-